OpenStudy (help!!!!):

11. Jim has 5 pieces of string. He needs to choose the piece that will be able to go around his 36-inch waist. His belt broke, and his pants are falling down. The piece needs to be at least 4 inches longer than his waist so he can tie a knot in it, but it cannot be more that 6 inches longer so that the ends will not show from under his shirt. Which of the following pieces of string will work the best? A. 3 feet B. 3 ¾ feet C. 3 ½ feet D. 3 ¼ feet E. 2 ½ feet

OpenStudy (mathstudent55):

36 inch waist must be at least 36 + 4 inches = 40 inches cannot be more than 36 + 6 inches = 42 inches to convert inches to feet, divide inches by 12 40 inches = 3 1/3 ft 42 inches = 3 1/2 ft The string must be more than 3 and one third foot but connot be longer than 3 and a half feet.

OpenStudy (anonymous):

Poor Jim! I know the feeling! :-) This is how I would approach the problem. He has a 36 inch waist and the belt needs to be 4 inches longer, so that's 40 inches. It can't be 6 inches longer, which is 42 inches. So it needs to be between 40 and 42 inches long. A foot is 12 inches, so 40 inches is three feet and 4 inches (3 and one-third feet) and 42 inches is three feet and 6 inches (3 and one-half feet). I'd say the answer is "C"! Hope that's helpful (this is my first day on the site!) If I'm right, we've helped Jim too!

OpenStudy (help!!!!):

i dont know which one to choose for best answer..........

OpenStudy (mathstudent55):

There is only one answer that is at least 3 1/3 ft long and at most 3 1/2 ft long.

OpenStudy (help!!!!):

well i was talking about the medal soo.... anyways thanks for the help

OpenStudy (mathstudent55):

Look at each answer: A. 3 ft, too short, bec it's less than 3 1/3 ft B. 3 3/4 ft, too long bec it's longer than 3 1/2 ft D. 3 1/4 ft, too short bec it's shorter than 3 1/3 ft E. 2 1/2 ft, too short bec it's shorter than 3 1/3 ft C 3 1/2 ft, exactly the logest length allowed. This is it